Calculate the units of unsaturation present in compounds with the following molecular formulas. Draw a structure consistent with the molecular formula and your calculated units of unsaturation.

a) C6H10O

b) C7H11N0

c) C2H2BrN

d) C10H7CIO
Answer #1

units of unsaturation - 26+2 +N-H-X (0.0.1) Creosbong, Hohydro sens, Na nitrogens X-halogeng. @ GHIDO U.o.U - 2x6+2+0-10-0 2present 2 means iring titibond 2 , zring present 211 2 IT-bond » so, the structure of Iringt i bond is. V.O.U = 2x7+ - b) e H- C H₂ Brn V.O. U = 272+2+1-2-13 2means an.bonds Bine en d) Golyclo . 0.0.0 = 2x16 +84 - - =7 I mean aring + 5 7 boods